OpenStudy (spareb665):

Express log[5](w/8) as a difference.

OpenStudy (kamille):

\[\log _{5}\frac{ w }{ 8 }\] did you wanted to write this?

OpenStudy (spareb665):

Yes

OpenStudy (kamille):

So, there is a formula \[\log _{a}b - \log _{a}c=\log _{a}\frac{ b }{ ? }\] Try to do it with your own numbers

OpenStudy (kamille):

where is "?" should be "c"

OpenStudy (kamille):

ar according to your problem: \[\log _{a}\frac{ b }{c}=\log _{a}b - \log _{a}\]

OpenStudy (spareb665):

log5 w - log5 8?

OpenStudy (kamille):

yes:)